OpenStudy (anonymous):

log2x+log2(x+3)−log2(2x−5)=3

jhonyy9 (jhonyy9):

log2(x(x+3))/(2x-5)=3log2(2)

OpenStudy (anonymous):

thanks, i got two answers 5 and 8

jhonyy9 (jhonyy9):

x2+3x=8(2x-5) x2+3x=16x-40 x2 -13x +40=0

OpenStudy (anonymous):

yeah that's it.thanks

jhonyy9 (jhonyy9):

(x-5)(x-8)